11.3.77 Q21 In​ November, the rain in a certain valley tends to fall in storms of several​ days' duration. The unconditional probability of rain on any given day of the month is 0.700. But the probability of rain on a day that follows a rainy day is 0.900​, and the probability of rain on a day following a non rainy day is 0.400. Find the probability of rain on November 1 and​ 2, but not on November 3.

Find the probability of rain on November 1 and​ 2, but not on November 3 is
Answer #1

a)

probability of rain on November 1 = 0.7( as no record of previous)

probability of rain on November 2 (if rain on nov 1) = 0.9

probability of not rain on November 3 (if rain on nov 2) = 1- 0.9 =0.1

probability of rain on November 1 and​ 2, but not on November 3 =

0.7 * 0.9 * 0.1 =0.063

Please revert back in case of any doubt.

Please upvote. Thanks in advance.